# Emergence Learning: A Rising Direction from Emergent Abilities and a Monosemanticity-Based Study

### 17 Dec 2023 v1

@@ 3.1.1 Metric Definition
Given a neuron $h$, we denote its historical samples under $m$ inputs as $\{h_{(1)},h_{(2)},\cdots,h_{(m)}\}$ and new value under the $(m+1)^{th}$ input as $h_{m+1}$. The Monosemantic Scale is defined as
$\phi(h_{(m+1)})=\frac{(h_{(m+1)}-\bar{h})^{2}}{S^{2}},$ (1)
@@
3.1.2 Metric Analysis
In this subsection, we state the effectiveness of our metric. Let’s begin by discussing two observations of monosemantic neurons.
  • High deviation of activation value. Qualitatively, a neuron is considered “activated” when its value for the current input deviates from the mean. For a monosemantic neuron, its value distribution is more skewed and incurs larger deviation. However, strictly defining “monosemantic” and “activated” is still under discussion in quantitative analysis (Elhage et al., 2022).
  • Low frequency of activation. Each monosemantic neuron only activates when its corresponding feature is inputted, which rarely happens considering the current datasets with steadily growing scales of samples and feature types.
The high deviation of the activation value ensures that the term $(h_{(m+1)}-\bar{h})^{2}$ in $\phi$ can effectively filter neurons with high monosemanticity. The low frequency of activation ensures that deviated values, when the neuron activates, hardly influence the value $\bar{h}$. $\bar{h}$ is mainly determined by the values when the neuron is deactivated.
On the other hand, our target only requires finding neurons that are monosemantic, whereas existing works need to first find the neuron-feature relationships (Gurnee et al., 2023; Trenton et al., 2023). Our solution focuses on handling monosemanticity and relaxes the requirement to find their corresponding features, eliminating the need for a predefined feature dataset.
@@ 3.1.2 Metric Analysis
Figure 2: Illustration of problems and solutions to inhibit monosemanticity. (a) A monosemantic neuron $h$ only activates (orange) for the feature “cat” with a high mean value ($=7$). $h$ is deactivated (blue) for other inputs with a small mean value ($\bar{h}=1$). (b) The first goal is to optimize the frontal model $f_{1}$ so that $h$ is less activated given the input “cat”. (c) The second goal is to optimize the followed model $f_{2}$ so that a correct output for “cat” does not solely rely on $h$. (d) Naive solution that fixes $h$ to a constant 1 without gradient. (e) Naive solution that decreases $h$ to $\bar{h}$ with a constant 6 without gradient. (f) Reverse Deactivation that first reverses $h$ then pushes the output value to $\bar{h}$ by adding a constant 8 without gradient. (g) All the methods can achieve the second goal by outputting $\bar{h}$ to $f_{2}$. As $\bar{h}$ provides little information, $f_{2}$ must learn to rely on other neurons. (h) When calculating the gradient, $f_{1}$ will find that $h$ is too small and tends to increase it (e.g., from 1 to 2). (i) Naive method 1 cannot be updated without gradient. Naive method 2 further increases the underlying $h$ activation (7 to 8). (j) Reverse Deactivation inherently deactivates $h$ (from 7 to 6). When a new batch arrives, $h$ activates less for “cat”.
3.1.3 Metric Computing
@@
Computing metric $\phi$ online is nontrivial due to the involvement of historical samples. Here we show that for each neuron, our proposed MS $\phi$ can be obtained by keeping track of 2 variables in constant time ($O(1)$). Since inputs are typically received in batches (with the number of new samples being greater than 1) during the training of deep learning models, we present the following lemma for training with a batch size of $b$.
Lemma 3.2.
@@
Denote $\mu_{m}$ as the value of the sample mean $\bar{h}$ given $m$ samples, while $\upsilon_{m}$ as the sample variance $S^{2}$. When the $(m+1)^{th}\sim(m+b)^{th}$ samples $h_{(m+1)}\sim h_{(m+b)}$ come, one can obtain the updated values via:
$\mu_{m+b}$ $=\frac{m\mu_{m}+b\mu_{b}^{\prime}}{m+b},$ (2)
$\upsilon_{m+b}$ $=\frac{mb(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)}+\frac{b\upsilon^{\prime}_{b}+(m-1)\upsilon_{m}}{m+b-1},$ (3)
@@
where $\mu_{b}^{\prime}=\frac{\sum_{i=1}^{b}h_{(m+i)}}{b}$ and $\upsilon^{\prime}_{b}=\frac{\sum_{i=1}^{b}(h_{(m+i)}-\mu^{\prime}_{b})^{2}}{b}$, which is of $O(1)$ time and memory complexity as $b$ is a constant.
The proof is given in Appendix. B and $\phi$ can be directly derived using Equation. (1).
@@
In the implementation, we introduce a forget mechanism as the model is updating, where the influence of previous samples should decay. For details, please refer to Algorithm. 1 in Appendix. D.
3.2 Inhibition of Monosemanticity
@@
Recall that given a monosemantic neuron $h$ and an input $X$ that contains its exclusive feature, we aim to optimize the model in two aspects: (1) decrease the activation level of $h$ when given $X$ (Figure. 2(b)); and (2) reduce the reliance of output $Z$ on the activation of $h$ (Figure. 2(c)).
In this subsection, we first demonstrate that directly deactivating monosemantic neurons can instead enhance monosemanticity. To address this unexpected phenomenon, we propose a method called Reversed Deactivation and provide the theoretical analysis.
3.2.1 Naive Deactivation
Without loss of generality, let’s focus on a single layer of neurons $H$ and a single neuron $h\in H$ that exhibits high monosemanticity. As defined in subsection. 3.1, such a neuron has a large $\phi(h)=(h-\bar{h})^{2}/S^{2}$, which is expected to be inhibited during model optimization. We can divide the optimization target into two parts: the frontal model $f_{1}$ before the layer, and the followed model $f_{2}$ after the layer.
@@ 3.2.1 Naive Deactivation
From the perspective of information theory, suppose that the distribution of neuron values follows a normal distribution, a neuron provides the least amount of information when its value equals its statistical mean ($\min_{h}[{I(h)}]=\min_{h}[{-\log(P(h))}]=-\log(P(\bar{h}))$), which is also its most inactive state. Thus, one can find two naive solutions to deactivate a neuron by replacing the original $h$ with $h^{\prime}$:
$(a)h^{\prime}=$ $\bar{h}_{ng}$ (4)
$(b)h^{\prime}=$ $h+(\bar{h}-h)_{ng},$ (5)
where subscript $\cdot_{ng}$ refers to “no gradient” scalar without trainable parameter. Two examples are given in Figure. 2(d,e)
@@
Both solutions ensure that $f_{2}$ receives an updated $H^{\prime}$, in which the value of $\mathcal{V}(h^{\prime})=\bar{h}$ provides little information. As $h^{\prime}$ is valueless, $f_{2}$ has to adjust its parameter to utilize information from other neurons in $H$ for good output $Z^{\prime}=f_{2}^{*}(H^{\prime},X)$. Such a process achieves our second goal of reducing the dependence of output $Z$ on the activation of $h$ (Figure. 2(g)).
@@
However, neither of the two solutions can achieve the first goal (i.e., training $f_{1}$ to inhibit the activation of $h$ for a single feature). To be more specific, method (a) wastes all the gradient for $f_{1}$ in generating $h$, preventing the related parameters from being updated. Method (b) outputs $\mathcal{V}(h^{\prime})=\bar{h}$ by compensating for the gap. As the model obtains a good result with $h$, by calculating the gradient, $f_{1}$ will be updated to push its output value from $\bar{h}$ to $h$, expressed as $\bar{h}+\delta(h-\bar{h})$, where $\delta>0$ up to the learning rate (Figure. 2(h)). Ironically, the actual output of $h$ is updated to $h+\delta(h-\bar{h})$, deviating from $\bar{h}$ by a factor of $1+\delta$:
$\frac{h+\delta(h-\bar{h})-\bar{h}}{h-\bar{h}}=1+\delta.$
@@
Therefore, these two simple solutions either contribute nothing to the deactivation of $h$ or even enhance its activation (Figure. 2(i)).
@@
(It is possible that previous researchers also recognized the negative effects of monosemanticity, but they may have discontinued their efforts after obtaining bad results from implementing these naive solutions.)
Figure 3: Overview of our method: (a) An arbitrary neural network framework. $X$ represents the input and $Z$ represents the output. $H_{i}$s represent hidden layers of neurons, and arrows indicate the dependency relationships. (b) Our module is inserted after $H_{3}$ and $H_{5}$, requiring no changes to the framework. (c) Details of our module applied to $H_{5}$. The input neurons are first analyzed using our metric. Once monosemantic neurons are identified, they are inhibited using RD. The resulting processed layer has the same shape as the input.
3.2.2 Reversed Deactivation
@@
To address the aforementioned problems, we propose our method called Reversed Deactivation (RD for short). Following the above-mentioned definitions, we replace the original $h$ with $h^{\prime}$ (Figure. 2(f)):
$h^{\prime}=-h+(\bar{h}+h)_{ng}.$ (6)
@@
Similar to baselines, RD also ensures the second goal of decreasing the dependence of output $Z$ on the activation of $h$. To be more specific, $f_{2}$ receives a value $\mathcal{V}(h^{\prime})=\mathcal{V}({-h+(\overline{h}+h)_{ng}})=\overline{h}$, which provides little information and requires $f_{2}$ to learn the information from other neurons in $H$ (Figure. 2(g)).
@@
Besides, RD can inhibit the activation level of $h$ when given $X$. In short, after calculating the gradient, $f_{1}$ will update the trainable parameter to push its output value from $\bar{h}$ to $h$ (Figure. 2(h)). Different from method (b), the gradient on $h$ is reversed. To achieve the same shifted value $\bar{h}+\delta(h-\bar{h})$ mentioned above, an insight into the update can be expressed as:
$\mathcal{V}(-\boldsymbol{h}+(\bar{h}+h)_{ng})$ $=\bar{h}$ (7)
$\rightarrow\mathcal{V}(-\boldsymbol{(h-\delta(h-\bar{h}))}+(\bar{h}+h)_{ng})$ $=\bar{h}+\delta(h-\bar{h})$ (8)
@@
The output of $f_{1}$ without post deactivation (i.e., $h-\delta(h-\bar{h})$) is closer to $\bar{h}$ with a factor $1-\delta$:
$\frac{h-\delta(h-\bar{h})-\bar{h}}{h-\bar{h}}=1-\delta,$
@@
Lemma 3.3.
Given a trained model $\mathcal{F}$ with $2$ continuous derivatives and a Lipschitz continuous gradient, where $\mathcal{F}$ achieves a desired output $Z$ with minimal loss $\mathcal{L}(Z)$, in which $Z=\mathcal{F}(X)=f_{2}(f_{1}(X),X)=f_{2}(H,X)$ for input $X$ based on its monosemantic neuron $h$ in layer $H$, suppose that $\mathcal{L}(f_{2}(\cdot))$ monotonically increases with $|h^{\prime}-h|$ for any other value $h^{\prime}$ that replaces $h$. Then, with a sufficiently small learning rate $l$, by updating the model $\mathcal{F}$ with gradient descent based on the neuron processed by the RD method, the activation of $h$ on input $X$ can be inhibited.
The proof is given in Appendix. C
3.3 Flexible Plug-in Module
@@
The details of MEmeL are displayed in Figure. 3. We present a general framework of a neural network in Figure. 3(a). The output $Z$ is derived from $X$ based on the hidden states $H_{i}$. Yellow arrows indicate the reliance of neuron layers on each other. Our method can be applied to any layer of neurons, such as $H_{3}$ and $H_{5}$ in Figure. 3(b), where the original neurons are adjusted by our modules to inhibit monosemanticity.
Taking $H_{5}$ as an example (Figure. 3(c)), we first detect monosemantic neurons using our MS metric, as introduced in subsection. 2.2. After identifying these neurons (colored red), we apply our Reverse Deactivation method, as described in subsection. 2.3, to each of them.
For adaptability, our module (i) does not require a specific input format and (ii) outputs $H_{5}^{\prime}$ in the same shape as the input $H_{5}$, ensuring format consistency during data propagation. Thus, no adjustments to the framework are needed to apply our module.
For lightweightness, neither of our two methods introduces any trainable parameters. Additionally, MEmeL focuses on presenting the idea of functionality induction. The methods we propose do not directly decrease monosemanticity during training, but instead induce the model to do so, supported by theoretical analysis. Once we have a well-trained model, the performance is expected to be robust even without induction. Therefore, during testing, the module can be directly removed without incurring any inference overhead.
The detailed algorithm is provided in the Appendix. E. We also apply two simple tricks for implementation: late start and variance compensation, to avoid unstable value fluctuations during the cold start.
Last but not least, we emphasize the proposition of the pipeline for emergent learning. Researchers can focus on separable directions: (1) improving the metric to better detect factors related to emergence, and (2) designing factor-oriented solutions for better performance.
4 Experiments
4.1 Setup
Table 1: Results on GLUE Test datasets, which are evaluated online (https://gluebenchmark.com/leaderboard). We follow the setting of BERT to demonstrate results on 8 datasets and calculate the average score. The scores are F1 scores for QQP and MRPC, Spearman correlations for STS-B, and accuracy scores for the other tasks. All metrics are the larger the better with best results in bold font.
@@ 4.1 Setup
BERTBASE 84.6/83.4 71.2 90.5 93.5 52.1 85.8 88.9 66.4 79.6
BERTMEmeL 84.9/83.8 71.6 91.3 93.7 53.8 85.9 88.9 67.8 80.2
Table 2: Results on ImageNet-1k dataset, where 3 sizes of Swin-Transformer pretrained on ImageNet-22k are used as backbones. The rows labeled “Original” and “MEmeL” refer to the results of models without and with MEemL, respectively. The metric used is top-1 accuracy, where a higher value indicates better performance. The best results are indicated in bold font.
Model Swin-T Swin-S Swin-B

### 15 Feb 2024 v2

@@ 3.1. Metric for Monosemanticity
Given a neuron $z_{i}\in\mathbf{z}$, we denote its historical samples under $m$ inputs $\{\mathbf{x}^{[1]},\mathbf{x}^{[2]},\cdots,\mathbf{x}^{[m]}\}$ as $\{z^{[1]}_{i},z^{[2]}_{i},\cdots,z^{[m]}_{i}\}$ and new value under the $(m+1)$-th input $\mathbf{x}^{[m+1]}$ as $z^{[m+1]}_{i}$. The Monosemantic Scale is defined as:
(7) $\phi(z^{[m+1]}_{i})=\frac{(z^{[m+1]}_{i}-\bar{z}_{i})^{2}}{S^{2}},$
@@
In the following contents, we will focus on this single neuron and use $z$ for simplicity.
As shown in Eq. (7), the measurement $\phi(\cdot)$ is proportional to the degree of deviation from the mean. The high deviation of the activation value ensures that the term $(z^{[m+1]}-\bar{z})^{2}$ in $\phi$ can effectively identify neurons with high monosemanticity. Besides, the metric is inversely proportional to the degree of fluctuation because the size of the deviation is also highly correlated with the fluctuation of the activation value on the current data set. Since $\bar{z}_{i}$ is mainly determined by the values when the neuron is deactivated, using the mean as the benchmark for evaluating deviations ensures that the defined neurons with high activation values are rare, i.e., low frequency.
As discussed in Sec. 1 and 2, prior works need to first find the neuron-feature relationships (Gurnee et al., 2023; Trenton et al., 2023) under the manually defined feature data set. Instead, we relax the requirement of discovering corresponding features and eliminate the need for a predefined feature set since we focus on finding monosemantic neurons.
@@
Figure 2. Illustration of problems and solutions to inhibit monosemanticity. (a) A monosemantic neuron $z$ only activates (orange) for the feature “cat” with a high mean value ($=7$). $z$ is deactivated (blue) for other inputs with a small mean value ($\bar{z}=1$). (b) The first goal is to optimize the frontal model $f_{1}$ so that $z$ is less activated given the input “cat”. (c) The second goal is to optimize the followed model $f_{2}$ so that a correct output for “cat” does not solely rely on $z$. (d) Zoom in on the original model at neuron $z$. (e) Naive solution that sets $z^{\prime}$ to a constant 1 without gradient. (f) Naive solution that decreases the value of $z$ to $\bar{z}$ with a constant 6 without gradient. (g) Reverse Deactivation that first reverses $z$ then pushes the output value to $\bar{z}$ by adding a constant 8 without gradient. (h) All the methods can achieve the second goal by outputting a value $\mathcal{V}(z^{\prime})=\bar{z}$ to $f_{2}$. As $\bar{z}$ provides little information, $f_{2}$ must learn to rely on other neurons. (i) When calculating the gradient, $f_{1}$ will find that $z$ is too small and tends to increase it (e.g., from 1 to 2). (j) Naive method (a) cannot update related parameters without gradient. (k) Naive method (b) further increases the underlying $z$ activation (7 to 8). (l) Reverse Deactivation inherently deactivates $z$ (from 7 to 6). When a new batch arrives, the updated $z^{*}$ activates less (=6) for “cat” compared with $z$.
@@
Metric Online Computing Guarantee. It is mandatory to compute the metric $\phi(\cdot)$ to avoid excessive computational burden caused by detecting monosemanticity. As discussed in Sec. 2, pioneer works (e.g., pairwise comparison in Eq. (6)) and other detection methods like probes necessitate training and offline inference for statistical analysis may be costly for online training. Here we show that for each neuron, our proposed MS $\phi(\cdot)$ can be obtained by keeping track of 2 variables in constant time ($O(1)$). Since inputs are typically received in batches (with the number of new samples being greater than 1) during the training of deep learning models, we present the following lemma for training with a batch size of $b$.
Lemma 3.2 ().
@@
Denote $\mu_{m}$ as the value of the sample mean $\bar{z}$ given $m$ samples, while $\upsilon_{m}$ as the sample variance $S^{2}$. When the $(m+1)^{th}\sim(m+b)^{th}$ samples ${z^{[m+1]},\cdots,z^{[m+b]}}$ come, one can obtain the updated values via:
(8) $\mu_{m+b}$ $=\frac{m\mu_{m}+b\mu_{b}^{\prime}}{m+b},$
(9) $\upsilon_{m+b}$ $=\frac{mb(\mu_{m}-\mu_{b}^{\prime})^{2}}{(m+b-1)(m+b)}+\frac{b\upsilon^{\prime}_{b}+(m-1)\upsilon_{m}}{m+b-1},$
@@
where $\mu_{b}^{\prime}=\frac{\sum_{i=1}^{b}z_{[m+i]}}{b}$ and $\upsilon^{\prime}_{b}=\frac{\sum_{i=1}^{b}(z_{[m+i]}-\mu^{\prime}_{b})^{2}}{b}$, which is of $O(1)$ time and memory complexity as $b$ is a constant.
@@
The proof is given in Appendix. C.1. In the implementation, we introduce a forgetting mechanism during model updates, where the influence of previous samples should decay. For details, please refer to Algorithm. 1 in Appendix. E.2.
3.2. Inhibition of Monosemanticity
After defining a quantitative metric in Sec. 3.1, it is intuitive to inhibit those monosemantic neurons directly. Unfortunately, simply deactivating monosemantic neurons will intensify the monosemanticity of neural networks. In this subsection, we will first prove the weakness of the native solution in Sec. 3.2.1. To address this unexpected phenomenon, we propose a theoretically supported reverse deactivation method in Sec. 3.2.2.
3.2.1. Naive Deactivation
@@
Recall that given a monosemantic neuron $z$ and an input $\mathbf{x}$ that contains its exclusive feature, we aim to optimize the model in two aspects: (1) decrease the activation level of $z$ when given $\mathbf{x}$ (Figure. 2(b)); and (2) reduce the reliance of output $\mathbf{o}$ on the activation of $z$ (Figure. 2(c)).
@@
Without loss of generality, we focus on the most monosemantic single neuron $z$ in a layer of neurons $\mathbf{z}$. As defined in Sec. 3.1, such a neuron has a large $\phi(z)=(z-\bar{z})^{2}/S^{2}$, which is expected to be inhibited during model optimization. From the perspective of information theory, suppose that the distribution of activation values follows a normal distribution, a neuron provides the least amount of information $I(z)$ when its value equals its statistical mean:
$\min_{z}[{I(z)}]=\min_{z}[{-\log(P(z))}]=-\log(P(\bar{z})),$
which is also its most inactive state. $P(\cdot)$ represents the probability density function of values of $z$.
Thus, a straightforward idea to deactivate $z$ is to modify its value to $\bar{z}$. We denote the modified neuron as $z^{\prime}$. One can find two naive solutions to deactivate a neuron:
(10) $\text{way}(a):\ \ z^{\prime}=$ $\bar{z}_{ng},$
(11) $\text{way}(b):\ \ z^{\prime}=$ $z+(\bar{z}-z)_{ng},$
@@
where subscript $\cdot_{ng}$ refers to a “no gradient” fixed-value scalar without trainable parameters, compared with $z\in\mathbf{z}=f_{1}(\mathbf{x})$ which is updatable. Two examples are given in Figure. 2(e,f). Both solutions ensure that $f_{2}$ receives an updated $z^{\prime}$, in which the value of $\mathcal{V}(z^{\prime})=\bar{z}$ provides little information. As $z^{\prime}$ is valueless, $f_{2}$ has to adjust its parameter to utilize information from other neurons in $\mathbf{z}$ for good output $\mathbf{o}^{*}=f_{2}^{*}(\mathbf{z}^{\prime},\mathbf{x})$. Such a process achieves our second goal of reducing the dependence of output $\mathbf{o}$ on the activation of $z$ (Figure. 2(l)).
@@
However, neither of the two solutions can achieve the first goal (i.e., training $f_{1}$ to inhibit the activation of $z$ for a single feature). To be more specific, method (a) wastes all the gradient for $f_{1}$ in generating $z$, preventing the related parameters from being updated. Method (b) outputs $\mathcal{V}(z^{\prime})=\bar{z}$ by compensating for the gap. As the model obtains a good result with $z$, by calculating the gradient, $f_{1}$ will be updated to push its output value from $\bar{z}$ to $z$, expressed as $\bar{z}+\delta(z-\bar{z})$, where $\delta>0$ up to the learning rate (Figure. 2(i)). Ironically, the actual output of $z$ is updated to $z+\delta(z-\bar{z})$, deviating from $\bar{z}$ by a factor of $1+\delta$:
$\frac{z+\delta(z-\bar{z})-\bar{z}}{z-\bar{z}}=1+\delta.$
@@
Therefore, these two simple solutions either contribute nothing to the deactivation of $z$ or even enhance its activation (Figure. 2(gk)).
@@
After delving deeper into the literature, it is possible that previous researchers also recognized the negative effects of monosemanticity, but they may have discontinued their efforts after obtaining bad results from implementing these naive solutions.
3.2.2. Reversed Deactivation
@@
To address the aforementioned problems, we propose our method called Reversed Deactivation (RD for short). Following the above-mentioned definitions, we replace the original $z$ with $z^{\prime}$ (Figure. 2(f)):
(12) $z^{\prime}=-z+(\bar{z}+z)_{ng}.$
@@
Similar to baselines, RD also ensures the second goal of decreasing the dependence of output $\mathbf{o}$ on the activation of $z$. To be more specific, $f_{2}$ receives a value $\mathcal{V}(z^{\prime})=\mathcal{V}({-z+(\overline{z}+z)_{ng}})=\overline{z}$, which provides little information and requires $f_{2}$ to learn the information from other neurons in $\mathbf{z}$ (Figure. 2(h)).
@@
Besides, RD can inhibit the activation level of $z$ when given $\mathbf{x}$. In short, after calculating the gradient, $f_{1}$ will update the trainable parameter to push its output value from $\bar{z}$ to $z$ (Figure. 2(i)). Different from method (b), the gradient on $z$ is reversed. To achieve the same shifted value $\bar{z}+\delta(z-\bar{z})$ mentioned above, an insight into the update can be expressed as:
(13) $\mathcal{V}(-{z}+(\bar{z}+z)_{ng})$ $=\bar{z}$
(14) $\rightarrow\mathcal{V}(-{(z-\delta(z-\bar{z}))}+(\bar{z}+z)_{ng})$ $=\bar{z}+\delta(z-\bar{z})$
@@
The output of $f_{1}$ without post deactivation (i.e., $z-\delta(z-\bar{z})$) is closer to $\bar{z}$ with a factor $1-\delta$:
$\frac{z-\delta(z-\bar{z})-\bar{z}}{z-\bar{z}}=1-\delta,$
@@
Lemma 3.3 ().
Given a trained model $f$ with $2$ continuous derivatives and a Lipschitz continuous gradient, where $f$ achieves a desired output $\mathbf{o}$ with minimal loss $\mathcal{L}(\mathbf{o})$, in which $\textbf{o}=f(\mathbf{x})=f_{2}(f_{1}(\mathbf{x}),\mathbf{x})=f_{2}(\mathbf{z},\mathbf{x})$ for input $\mathbf{x}$ based on its monosemantic neuron $z$ in layer $\mathbf{z}$, suppose that $\mathcal{L}(f_{2}(\cdot))$ monotonically increases with $|z^{\prime}-z|$ for any other value $z^{\prime}$ that replaces $z$. Then, with a sufficiently small learning rate $l$, by updating the model $f$ with gradient descent based on the neuron processed by the RD method, the activation of $z$ on input $\mathbf{x}$ can be inhibited.
The proof is given in Appendix. C.2.
Figure 3. Overview of our method: (a) An arbitrary neural network framework. $\mathbf{x}$ represents the input and $\mathbf{o}$ represents the output. $\mathbf{z}$s represent hidden layers of neurons, and arrows indicate the dependency relationships. (b) Our module is inserted after $\mathbf{z}^{3}$ and $\mathbf{z}^{5}$, requiring no changes to the framework. (c) Details of our module applied to $\mathbf{z}^{5}$. The input neurons are first analyzed using our metric. Once monosemantic neurons are identified, they are inhibited using RD. The resulting processed layer has the same shape as the input.
Additionally, we conduct validation experiments to compare the optimization outputs based on different methods in subsection. 4.3. The results are consistent with the theory and the analysis of naive methods and RD, showing that RD is effective in inhibiting monosemanticity.
3.3. Flexible Plug-in Module
@@
The details of MEmeL are displayed in Figure. 3. We present a general framework of a neural network in Figure. 3(a). The output $\mathbf{o}$ is derived from $\mathbf{x}$ based on the layers of neurons $\{\mathbf{z}\}$. Yellow arrows indicate the reliance of neuron layers on each other. Our method can be applied to any layer of neurons, such as $\mathbf{z}^{3}$ and $\mathbf{z}^{5}$ in Figure. 3(b), where the original neurons are adjusted by our modules to inhibit monosemanticity.
Taking $\mathbf{z}^{5}$ as an example (Figure. 3(c)), we first detect monosemantic neurons using our MS metric, as introduced in subsection. 3.1. After identifying these neurons (colored red), we apply our Reverse Deactivation method, as described in subsection. 3.2, to each of them.
For adaptability, our module (i) does not require a specific input format and (ii) outputs $(\mathbf{z}^{5})^{\prime}$ in the same shape as the input $\mathbf{z}^{5}$, ensuring format consistency during data propagation. Thus, no adjustments to the framework are needed to apply our module.
For lightweightness, neither of our two methods introduces any trainable parameters. Additionally, MEmeL focuses on presenting the idea of functionality induction. The methods we propose do not directly decrease monosemanticity during training, but instead induce the model to do so, supported by theoretical analysis. Once we have a well-trained model, the performance is expected to be robust even without induction. Therefore, during testing, the module can be directly removed without incurring any inference overhead. The results of the validation experiments also support our analysis of removing MEmeL during testing. (see Appx. F.1).
Table 1. Results on GLUE Test datasets. We follow the setting of BERT to demonstrate results on 8 datasets and calculate the average score. The scores are F1 scores for QQP and MRPC, Spearman correlations for STS-B, and accuracy scores for the other tasks. All metrics are the larger the better with best results in bold font.
@@
Original 84.6/83.4 71.2 90.5 93.5 52.1 85.8 88.9 66.4 79.6
Naive (a) 84.3/83.6 71.7 90.6 93.8 52.1 85.8 88.2 66.4 79.6
Naive (b) 84.7/84.1 71.6 90.6 93.6 51.8 86.5 87.2 68.0 79.8
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 66.4 80.0
MEmeL-Tune 84.8/83.9 71.7 91.2 93.7 55.7 86.6 89.0 68.1 80.5
The detailed algorithm is provided in the Appendix. D.2. We also apply two simple tricks for implementation: late start and variance compensation, to avoid unstable value fluctuations during the cold start.
Last but not least, we emphasize the proposition of the pipeline for emergence-based learning. Researchers can focus on separable directions: (1) improving the metric to better detect factors related to scale change, and (2) designing factor-oriented solutions for better performance.
4. Experiments
In this section, we first introduce the basic experimental setup in Sec. 4.1, then present the main empirical studies in Sec. 4.2. We show a case study that the proposed Reverse Deactivation is more powerful in inhibition monosematicity than the naive methods in Sec. 4.3. Furthermore, we discuss the potentials and limitations of MEmeL in Sec. 4.4. Note the experiment about the impact of removing MEmeL during test is provided in Appx. F.1 due to the limited space.
4.1. Experimental Setup
Table 2. The experimental results of Swin-Transformer on the ImageNet data and ConvGRU on the data HKO-7. For results on ImageNet-1k dataset, 3 Swin-Transformers pretrained on ImageNet-22k are used as backbones. The metric used is top-1 accuracy, where a higher value indicates better performance. For results on HKO-7 dataset, we initially train a ConvGRU model for 20k steps to create the base model. The metrics used are B-MSE and B-MAE, where a smaller value indicates better performance. The best results are in bold fonts.
Model Swin-T Swin-S Swin-B B-MAE B-MSE